OpenStudy (idealist10):

A 96 lb weight is dropped from rest in a medium that exerts a resistive force with magnitude proportional to the speed. Find its velocity as a function of time if its terminal velocity is -128 ft/s.

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

 The total force acting on the object is the sum of the resisting forces upward and the downward gravitational force. \[m\frac{dv}{dt}=mg-kv\] Solve this ODE for \(v\) (a function of \(t\)). The terminal velocity is the velocity attained as \(t\to\infty\), but in this case you want to be able to use it as a sort of initial condition. At terminal velocity, acceleration is zero; this is because the velocity attains a constant maximum. So, the terminal velocity must occur when \(a=\dfrac{dv}{dt}=0\), i.e. when \(mg-kv=0\), or \(v=\dfrac{mg}{k}\).